The width of a rectangle is 2 yards shorter than its length. the area of the rectangle is 48 square yards. find the length and width of the rectangle.
Answer by TimothyLamb(4379)

w = L - 2
Lw = 48
---
Lw = 48
L(L - 2) = 48
LL - 2L = 48
L^2 - 2L - 48 = 0
---
the above quadratic equation is in standard form, with a=1, b=-2, and c=-48
---
to solve the quadratic equation, plug this:
1 -2 -48
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the roots of the quadratic are:
h = 8
h = -6
---
the negative root doesn't make sense for a length, so use the positive root:
---
Answer:
L = 8 yards
w = 6 yards
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php